You are planning to make monthly deposits of $170 into a retirement account that pays 12 percent interest compounded monthly. If your first deposit will be made one month from now, how large will your retirement account be in 14 years?
Answer #1

[(1+г) — 1 FV of Annuity %3D Р т Р- Рeriodic Pаутent г3 rate per period п 3 питber o ff periods

For this question, n = 14 year * 12 = 168 months, r = 12%/12 = 1% (per month)

FV = 170 * [\frac{(1+0.01)^{168} -1}{0.01}]

FV = 170 * 432.097

FV = $73,456.49
Answer #2

SOLUTION :


Monthly compounding.


Monthly deposit, A = 170 ($)


Monthly interest rate, r = 12/12 =1% = 0.01 in decimals

=> 1 = r = 1.01


Periods, n = 14 * 12 = 168 months


So, 


FV after 14 years (168 months)


= A((1 + r)^n - 1) / r

= 170 * (1.01^168 - 1) / 0.01

= 73456.49 ($) (ANSWER).

answered by: Tulsiram Garg